OpenStudy (anonymous):

A manufacturer believes that the net profit P is related to the advertising expenditure x (both in thousands of dollars by P(x)=10+46x-1/2x^2. What advertising expenditure will maximize the profit?

OpenStudy (radar):

\[(1/2)x ^{2}\] or\[1/2x ^{2}\] which is it?

OpenStudy (radar):

going with the first dP/dx =-x+46 let =0, then x=46 (thousand dollars)

OpenStudy (anonymous):

Okay...sorry, slow pc

OpenStudy (anonymous):

P(x)=10+46x-1/2x^2 That is (-1/2x)^2

OpenStudy (radar):

OK thats what I thought, $46,000

OpenStudy (anonymous):

Thank you very much